OpenStudy (anonymous):

how to find the exact solution of: sin(4(x-pi/4))=0

OpenStudy (michele_laino):

hint: if: \[\sin \theta = 0\] then: \[\theta = k\,\pi ,\quad k = 0,\, \pm 1,\, \pm 2,...\]